# Improving Generalization of Neural Combinatorial Optimization for Vehicle Routing Problems via Test-Time Projection Learning

### v2

@@ 6.1 Experimental Results
We further test our method on TSPLIB [59] and CVRPLIB [60]. The detailed results are demonstrated in the Table 3. When using greedy construction for inference, LEHD purely with our designed projection can achieve a better result than other construction-based methods on all groups of instances. Using the MVDF and RRC further strengthens our performance, which yields a significant gap between the second-best methods.
@@ 6.2 Ablation Study
To validate the effectiveness of the LLM-designed projection strategy, we design two variants of our method: 1) w/o proj: LEHD only uses KNN to reduce the search space during inference. 2) seed proj: The projection strategy is replaced with the input seed strategy in [15]. As exhibited in the Table. 4, the projection strategy designed from the TTPL framework consistently outperforms the two variants on all scales.
@@
To assess the efficacy of our proposed MVDF, we compare our complete method against two configurations: 1) w/o MVDF: LEHD utilizes the designed projection but without MVDF during inference; 2) POMO aug: test instances are first augmented using the POMO augmentation [7], and then solved with LEHD equipped with the designed projection. The complete results are shown in Table 5. Our MVDF approach outperforms the w/o MVDF variant in 5 out of 5 cases and the POMO aug variant in 3 out of 5 cases. Notably, MVDF demonstrates particular effectiveness when solving large-scale instances.
@@ 6.3 Versatility study
The robustness of TTPL-MVDF is evaluated on cross-distribution instances (cluster, explosion, and implosion) from [15]. Compared against LEHD without projection and LEHD with seed projection, TTPL-MVDF demonstrates a significant improvement in the base model’s robustness. These findings (see Appendix D.2 for full results) validate the effectiveness of our proposed method in handling distributional shifts.
@@ A.1 Review for Non-Constructive Method
Despite direct learning to construct a VRP solution, other methods work closely with search and improvement methods [63, 64, 65]. Early studies [66] utilize a graph convolutional network to generate heatmaps indicating the probability of edges belonging to optimal solutions for TSP instances. Subsequently, an approximate solution can be obtained via beam search [66], Monte Carlo tree search [19], dynamic programming [67], and guided local search [68]. While initially developed for small-scale problems, recent studies extend heatmap-based methods to larger TSP instances [19, 69]. However, the adaptability of these methods can be limited by their reliance on carefully tuned search strategies. Furthermore, the heatmap representation primarily focuses on edge probabilities, which may restrict applicability to problems with complex constraints.
Another research direction involves learning-augmented algorithms, where neural networks enhance traditional heuristic solvers. This includes accelerating solvers by learning operation selection [63, 70, 71], guided improvement [72, 73, 74], and problem decomposition [45, 75]. Many improvement-based methods propose to iteratively refine feasible solutions [76, 77, 70]. Nonetheless, these methods typically depend on well-designed heuristic operators or advanced solvers [78, 57] for solving different problems.
@@ A.2 Review for LLM-AHD Method
Recent advancements demonstrate the significant potential of Large Language Models (LLMs) in automating the design of high-performing heuristics for complex computational problems. This methodology typically involves maintaining a population of elite heuristic functions, which are evaluated and ranked based on their fitness on a designated dataset. The LLM is then iteratively prompted with high-performing functions from this population to generate potentially superior heuristic candidates. EoH [52] and Funsearch [79] pioneer the application of LLM to design a population-based evolutionary computation procedure. Building on this, ReEvo [80] incorporates a reflection mechanism to augment the LLM’s reasoning capabilities during heuristic generation. Addressing the challenge of premature convergence, HSEvo [81] introduces diversity metrics and a harmony search algorithm to enhance population diversity without degrading solution quality. In a different approach, MCTS-AHD [53] conceptualizes the task as a search problem, modeling the heuristic design process as a Monte Carlo tree search to systematically explore the solution space.
@@ B.1 Prompts of TTPL
Our LLM-driven prompt framework for projection evolution consists of four core strategies (E1, E2, M1, M2) [52], each structured around three components: task description, strategy-specific instructions, parent projection, and template program. In the following, we detail their design principles:
@@
The prompt engineering for normalization evolution adopts a standardized template structure across strategies. The template code for TSP and CVRP builds on the INViT framework [15].
@@ B.2 EoH Algorithm
The EoH is a population-based framework for LLM-based AHD [52]. The algorithm optimizes heuristics by iteratively applying LLM-guided mutation, crossover, and selection operations, enabling the exploration of diverse heuristic designs.
@@ D.1 TSPLIB and CVRPLIB Results
We evaluate the proposed method in four benchmark scenarios: TSPLIB [59] and CVRPLib [60] instances with comparisons in both greedy and post-search frameworks. We compare our method against five baselines: INViT [15], SIGD [62], BQ [42], LEHD [18] and GLOP [13]. For TSPLIB, all experiments are conducted on large-scale instances ranging from 1,000 to 85,900 nodes to assess scalability and generalization.
@@ D.2 Result in Cross-Distributions
To evaluate robustness against distribution shifts, we tested TTPL on TSP and CVRP instances under three distinct spatial distributions (Clustered, Explosion, and Implosion)[15] on scales from 1,000 to 100,000 nodes.
@@ E.1 Adaptability of TTPL
To further validate the effectiveness of TTPL framework, we utilize our proposed framework on different base models. Specifically, we compare POMO [7], LEHD [18], BQ [42], and SIGD [62], four constructive models, with and without TTPL projection. The test instances are set to TSP1K/5K/10K with the same settings as the previous. The results are shown in the Table. 12
@@ F.2 Solution Visualizations of Large-scale TSPLIB Instances
Figure 7: The solution visualizations of TSPLIB[59] instances with different scales, the solutions are all generated by TTPL aug RRC1000.
@@ Appendix G License
LKH3[57] Code http://webhotel4.ruc.dk/keld/research/LKH-3/ Available for academic research use
HGS[58] Code https://github.com/chkwon/PyHygese MIT License
Concorde[82] Code https://github.com/jvkersch/pyconcorde BSD 3-Clause License
POMO[7] Code https://github.com/yd-kwon/POMO MIT License
LEHD[18] Code https://github.com/CIAM-Group/NCO_code/tree/main/ single_objective/LEHD MIT License
BQ[42] Code https://github.com/naver/bq-nco $C$ BY-NC-SA 4.0
GLOP[13] Code https://github.com/henry-yeh/GLOP MIT License
H-TSP[12] Code https://github.com/Learning4Optimization-HUST/H-TSP MIT License
DIFUSCO[28] Code https://github.com/Edward-Sun/DIFUSCO MIT License
INViT[15] Code https://github.com/Kasumigaoka-Utaha/INViT Available for academic research use
ELG[14] Code https://github.com/gaocrr/ELG MIT License
SIGD[62] Code https://github.com/grimmlab/gumbeldore Available for academic research use
EoH[52] Code https://github.com/FeiLiu36/EoH MIT License
LLM4AD[83] Code https://github.com/Optima-CityU/LLM4AD MIT License
TSPLIB[59] Dataset http://comopt.ifi.uni-heidelberg.de/software/TSPLIB95/ Available for any non-commercial use
CVRPLib[60] Dataset http://vrp.galgos.inf.puc-rio.br/index.php/en/ Available for academic research use

### v3

@@ 6.1 Experimental Results
We further test our method on TSPLIB [reinelt1991tsplib] and CVRPLIB [uchoa2017new]. The detailed results are demonstrated in the Table 3. When using greedy construction for inference, LEHD purely with our designed projection can achieve a better result than other construction-based methods on all groups of instances. Using the MVDF and RRC further strengthens our performance, which yields a significant gap between the second-best methods.
@@ 6.2 Ablation Study
w/o proj 24.2 (4.49%) 0.1s 57.5 (12.80%) 1.4s 84.3 (17.42%) 2.7s 236.7 (48.00%) 13.2s 500.4 (121.42%) 27.8s
seed proj 23.8 (2.96%) 0.1s 52.7 (3.32%) 1.5s 74.5 (3.77%) 2.9s 165.4 (3.42%) 14.4s 233.3 (3.24%) 29.6s
TTPL 23.73 (2.65%) 0.1s 52.63 (3.25%) 1.4s 74.39 (3.63%) 2.9s 165.19 (3.29%)) 14.3s 233.26 (3.22%) 28.9s
To validate the effectiveness of the LLM-designed projection strategy, we design two variants of our method: 1) w/o proj: LEHD only uses KNN to reduce the search space during inference. 2) seed proj: The projection strategy is replaced with the input seed strategy in [fang2024invit]. As exhibited in the Table. 4, the projection strategy designed from the TTPL framework consistently outperforms the two variants on all scales.
@@
To assess the efficacy of our proposed MVDF, we compare our complete method against two configurations: 1) w/o MVDF: LEHD utilizes the designed projection but without MVDF during inference; 2) POMO aug: test instances are first augmented using the POMO augmentation [kwon2020pomo], and then solved with LEHD equipped with the designed projection. The complete results are shown in Table 5. Our MVDF approach outperforms the w/o MVDF variant in 5 out of 5 cases and the POMO aug variant in 3 out of 5 cases. Notably, MVDF demonstrates particular effectiveness when solving large-scale instances.
@@ 6.3 Versatility study
The robustness of TTPL-MVDF is evaluated on cross-distribution instances (cluster, explosion, and implosion) from [fang2024invit]. Compared against LEHD without projection and LEHD with seed projection, TTPL-MVDF demonstrates a significant improvement in the base model’s robustness. These findings (see Appendix D.2 for full results) validate the effectiveness of our proposed method in handling distributional shifts.
@@ A.1 Review for Non-Constructive Method
Despite direct learning to construct a VRP solution, other methods work closely with search and improvement methods [chen2019learning, li2018combinatorial, hottung2021learning]. Early studies [joshi2019efficient] utilize a graph convolutional network to generate heatmaps indicating the probability of edges belonging to optimal solutions for TSP instances. Subsequently, an approximate solution can be obtained via beam search [joshi2019efficient], Monte Carlo tree search [fu2021generalize], dynamic programming [kool2022deep], and guided local search [hudson2021graph]. While initially developed for small-scale problems, recent studies extend heatmap-based methods to larger TSP instances [fu2021generalize, qiu2022dimes]. However, the adaptability of these methods can be limited by their reliance on carefully tuned search strategies. Furthermore, the heatmap representation primarily focuses on edge probabilities, which may restrict applicability to problems with complex constraints.
Another research direction involves learning-augmented algorithms, where neural networks enhance traditional heuristic solvers. This includes accelerating solvers by learning operation selection [chen2019learning, ma2021learning, d2020learning], guided improvement [cappart2021combining, hottung2020neural, xin2021neurolkh], and problem decomposition [li2021learning, song2020general]. Many improvement-based methods propose to iteratively refine feasible solutions [barrett2020exploratory, wu2021learning, ma2021learning]. Nonetheless, these methods typically depend on well-designed heuristic operators or advanced solvers [optimization2014inc, helsgaun2017extension] for solving different problems.
@@ A.2 Review for LLM-AHD Method
Recent advancements demonstrate the significant potential of Large Language Models (LLMs) in automating the design of high-performing heuristics for complex computational problems. This methodology typically involves maintaining a population of elite heuristic functions, which are evaluated and ranked based on their fitness on a designated dataset. The LLM is then iteratively prompted with high-performing functions from this population to generate potentially superior heuristic candidates. EoH [liu2024evolution] and Funsearch [romera2024mathematical] pioneer the application of LLM to design a population-based evolutionary computation procedure. Building on this, ReEvo [ye2024reevo] incorporates a reflection mechanism to augment the LLM’s reasoning capabilities during heuristic generation. Addressing the challenge of premature convergence, HSEvo [dat2025hsevo] introduces diversity metrics and a harmony search algorithm to enhance population diversity without degrading solution quality. In a different approach, MCTS-AHD [zheng2025monte] conceptualizes the task as a search problem, modeling the heuristic design process as a Monte Carlo tree search to systematically explore the solution space.
@@ B.1 Prompts of TTPL
Our LLM-driven prompt framework for projection evolution consists of four core strategies (E1, E2, M1, M2) [liu2024evolution], each structured around three components: task description, strategy-specific instructions, parent projection, and template program. In the following, we detail their design principles:
@@
The prompt engineering for normalization evolution adopts a standardized template structure across strategies. The template code for TSP and CVRP builds on the INViT framework [fang2024invit].
@@ B.2 EoH Algorithm
The EoH is a population-based framework for LLM-based AHD [liu2024evolution]. The algorithm optimizes heuristics by iteratively applying LLM-guided mutation, crossover, and selection operations, enabling the exploration of diverse heuristic designs.
@@ D.1 TSPLIB and CVRPLIB Results
We evaluate the proposed method in four benchmark scenarios: TSPLIB [reinelt1991tsplib] and CVRPLib [uchoa2017new] instances with comparisons in both greedy and post-search frameworks. We compare our method against five baselines: INViT [fang2024invit], SIGD [pirnay2024self], BQ [drakulic2023bq], LEHD [luo2023neural] and GLOP [ye2024glop]. For TSPLIB, all experiments are conducted on large-scale instances ranging from 1,000 to 85,900 nodes to assess scalability and generalization.
@@ D.2 Result in Cross-Distributions
To evaluate robustness against distribution shifts, we tested TTPL on TSP and CVRP instances under three distinct spatial distributions (Clustered, Explosion, and Implosion)[fang2024invit] on scales from 1,000 to 100,000 nodes.
@@ E.1 Adaptability of TTPL
To further validate the effectiveness of TTPL framework, we utilize our proposed framework on different base models. Specifically, we compare POMO [kwon2020pomo], LEHD [luo2023neural], BQ [drakulic2023bq], and SIGD [pirnay2024self], four constructive models, with and without TTPL projection. The test instances are set to TSP1K/5K/10K with the same settings as the previous. The results are shown in the Table. 12
@@ F.2 Solution Visualizations of Large-scale TSPLIB Instances
Figure 7: The solution visualizations of TSPLIB[reinelt1991tsplib] instances with different scales, the solutions are all generated by TTPL aug RRC1000.
@@ Appendix G License
LKH3[helsgaun2017extension] Code http://webhotel4.ruc.dk/keld/research/LKH-3/ Available for academic research use
HGS[vidal2022hybrid] Code https://github.com/chkwon/PyHygese MIT License
Concorde[applegate2006concorde] Code https://github.com/jvkersch/pyconcorde BSD 3-Clause License
POMO[kwon2020pomo] Code https://github.com/yd-kwon/POMO MIT License
LEHD[luo2023neural] Code https://github.com/CIAM-Group/NCO_code/tree/main/ single_objective/LEHD MIT License
BQ[drakulic2023bq] Code https://github.com/naver/bq-nco $C$ BY-NC-SA 4.0
GLOP[ye2024glop] Code https://github.com/henry-yeh/GLOP MIT License
H-TSP[pan2023h] Code https://github.com/Learning4Optimization-HUST/H-TSP MIT License
DIFUSCO[sun2023difusco] Code https://github.com/Edward-Sun/DIFUSCO MIT License
INViT[fang2024invit] Code https://github.com/Kasumigaoka-Utaha/INViT Available for academic research use
ELG[gao2023towards] Code https://github.com/gaocrr/ELG MIT License
SIGD[pirnay2024self] Code https://github.com/grimmlab/gumbeldore Available for academic research use
EoH[liu2024evolution] Code https://github.com/FeiLiu36/EoH MIT License
LLM4AD[liu2024llm4ad] Code https://github.com/Optima-CityU/LLM4AD MIT License
TSPLIB[reinelt1991tsplib] Dataset http://comopt.ifi.uni-heidelberg.de/software/TSPLIB95/ Available for any non-commercial use
CVRPLib[uchoa2017new] Dataset http://vrp.galgos.inf.puc-rio.br/index.php/en/ Available for academic research use